OpenStudy (anonymous):

Solve using elimination: -3x = 14 + 2y 3y − 8x = -21

OpenStudy (anonymous):

@MagicallyFarting

OpenStudy (mathstudent55):

Write the rwo equations as \(ax + by = c\) one above the other.

OpenStudy (mathstudent55):

Here's the first equation in that form: -3x - 2y = 14 Now do the same for the second equation and write it below the first one.

OpenStudy (anonymous):

is it be 8x-3y=-21

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Almost. You need to be careful with the signs. 3y − 8x = -21 is the same as 3y + (-8x) = -21 Notice that the -8x term is negative; it's really -8x, and the 3y term is positive, so it turns into -8x + 3y = -21

OpenStudy (mathstudent55):

Now we write the new equations one above the other and we are ready to start with the elimination method. -3x - 2y = 14 -8x + 3y = -21

OpenStudy (anonymous):

so what do i do now and i did :) write it in elimination method

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

The elimination method works by eliminating a variable. To do that, you need to add both equations and make sure that one variable adds to zero. The variables we have now, -3x and -8x add to -11x, and -2y and 3y add to y. They don't eliminate by addition, so we first need to modify them to be able to eliminate one variable by addition.

OpenStudy (anonymous):

that means its one whole equation like -11x+y=-21

OpenStudy (mathstudent55):

No. First, the numbers on the right would also have to be added, to give 14 + (-21) = -7, but it still wouldn't help becauyse you still have x and y in the equation. We need to eliminate one variable. This is how we do it. Notice the first equation has -2y, and the second equation has 3y. If we multiply the entire first equationn by 3, the y term will become -6y. If we multiply the entire second equation by 2, the y term will become 6y. -6y and 6y do add to zero, thus eliminating the y variable, so that's what we need to dodo.

OpenStudy (anonymous):

so i got 9x-6y=42 -16x+6y=-42

OpenStudy (mathstudent55):

Excellent. If you line them up, you'll be able to add more easily because the x's and the y's will line up: 9x - 6y = 42 -16x + 6y = -42 ---------------- add

OpenStudy (anonymous):

i got -7x+y=0

OpenStudy (texaschic101):

they y's cancelled out

OpenStudy (anonymous):

ok so it -7x=0

OpenStudy (mathstudent55):

Canceling out means no y left. You have a y there. Careful. We are trying to eliminate y. -6y + 6y = 0 9x - 6y = 42 -16x + 6y = -42 ---------------- add -7x = 0 See, no y variable left.

OpenStudy (mathstudent55):

You got it now. Now that we have no y, and a simple equation in x, you can solve -7x = 0 for x. Divide both sides by -7.

OpenStudy (texaschic101):

excellent explaining math student....I always have trouble explaining elimination problems to people

OpenStudy (mathstudent55):

@texaschic101 Thanks.

OpenStudy (anonymous):

x=0

OpenStudy (mathstudent55):

Great. Now that we have a value for x, insert that x value into x in one of the original equations.

OpenStudy (mathstudent55):

-3x = 14 + 2y -3(0) = 14 + 2y Now solve for y. First simplify the left side. Then subtract 14 from both sides.

OpenStudy (anonymous):

i got -7

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Excellent. You just solved a system of equations by the elimination method. Good job. The solution is: x = -7, y = 0

OpenStudy (anonymous):

awh thanks do you mind helping me on a few other questions also

OpenStudy (mathstudent55):

no problem. pls start a new post

OpenStudy (anonymous):

i tagged you in it